Please help me solve this linear system using the elimination method: 2x-3y=2
-5x+2y=-16
My work: 2x-3y=2 ..... 5(2x-3y=2)= 10x-15y=10
........-5x+2y=-16 ... 2(-5x+2y=-16)=-10x+4y=-32
...........................................-11y=-42...-11/-11= y...-42/-11=3/4
Answer I got: y=3/4 One Solution

I have tried multiplying it like this: 5(2x-3y=2) and 2(-5x+2y=-16) to get rid of x, then I subtracted: -15y=10 and 4y=-32 and got -11y=-42, and to get y by its self, i divided -11 by -42 and it left me with y=3/4 and it's a one solution
But I am not sure if this is correct.

Found 2 solutions by jim_thompson5910, richwmiller:
Answer by jim_thompson5910(35256)

You made a mistake when you added -32 to 10. You got the incorrect value of -42 when it's really -22


So it should be

-11y = -22

and solving for y gets you

y = 2

Now that you have this y value, use it to find x to complete the answer.

Answer by richwmiller(17219)

It is not correct.
x=4 y=2
You said:
2x-3y=2
5(2x-3y=2)= 10x-15y=10
5x+2y=-16 2(-5x+2y=-16)=-10x+4y=-32
add the two
-11y=-22 not -42
since 10-32=-22
y=2
2x-3*2=2
2x-6=2
2x=8
x=4